# Power allocation for cell-free MIMO integrated sensing and communication

### v1

@@ V-A Modified Conjugate Gradient Algorithm
where $\nabla^{2}L({\bm{\rho}}_{i})$ is the Hessian matrix of $L({\bm{\rho}})$ at ${\bm{\rho}}_{i}$. According to (39), we have ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\leq 0$. If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there exists a non-negative step size bound $\upsilon_{b}$ for given $\epsilon_{L}\in[0,1)$ such that, when $\upsilon_{i}\leq\upsilon_{b}$, we have
$f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (47)
When ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$, (47) always holds. In contrast, we have $\upsilon_{b}=\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}$ when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}>0$.
@@
$L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (48)
According to (39), the term on the right-hand side of (48) is always non-positive. Thus, (48) contributes to (38). The ILS controlling factor $\epsilon_{L}$ is used to control the decrement of the function $L({\bm{\rho}})$ by moving a step $\upsilon_{i}$. When (48) is unmet for given $\upsilon_{i}$, it is scaled by,
$\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$ (49)
where the scaling factor $\sigma_{\upsilon}\in(0,1)$ and an initial step size is denoted as $\upsilon_{i}=\upsilon_{0,i}$.
On the other hand, the penalty function does not inherently guarantee that the individual power budget constraints at each transmitter are satisfied during the optimization process. As a result, negative PA values may emerge during intermediate iterations. These infeasible PA values can lead to misleading CRLB evaluations, such as yielding negative CRLBs instead of infinite ones, thereby artificially satisfying the CRLB constraints. To prevent unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is essential to enforce non-negative PAs at each iteration step. To this end, the step size $\upsilon_{i}$ needs to satisfy,
$0<\upsilon_{i}\leq\upsilon_{i}^{\rm b}$ (50)
where $\upsilon_{i}^{\rm b}$ is a step size upper bound given by
@@
with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$. Thus, the initial step size for ILS can be selected as,
$\upsilon_{0,i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (52)
with $\epsilon_{\upsilon}\in(0,1]$. Note that there must exist ${d}_{n,i}<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$ due to $\sum_{n=1}^{N}{d}_{n,i}=0$. The ILS method characterized in (48)-(52) can be seen as the extension and application of the Armijo’s Rule [57, 56] in our work.
A termination condition is significant for the iterative algorithm. A natural choice, based on the ILS condition in (48), is that the absolute difference in objective function values between consecutive iterations satisfies $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$ where $\epsilon$ is a small positive threshold. To avoid numerical instability arising from the magnitude of the objective function itself, we apply a normalized termination condition, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, as employed in both the MCG-based and MSD-based algorithms.
We now present the Penalty-function and Projection-based Modified Conjugate Gradient algorithm with inexact line search (PP-MCG-ILS). A large penalty factor may lead to computational issues, such as an ill-conditioned Hessian matrix, which can adversely affect the convergence of the MCG algorithm. To mitigate this, we adopt a strategy of progressively increasing the penalty factor across iterations. For given penalty factor $\mu_{k}$, the MCG algorithm is executed until its termination condition is satisfied. Then, the penalty factor is updated as $\mu_{k+1}=\varphi\mu_{k}$ for the next loop of penalty-based iteration, where $\varphi>1$ is a scaling constant. This process continues until the convergence criterion $\mu_{k}\alpha(\bm{\rho})\leq\epsilon_{\mu}$ is met, with $\epsilon_{\mu}$ being a small, user-defined threshold [54]. This approach ensures numerical stability and gradual enforcement of the constraint. The complete PP-MCG-ILS algorithm is summarized in Algorithm 1.
@@
1:  Step size: Calculate $\upsilon_{i}^{\rm b}$ according to (51); set $\upsilon_{i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ , while $L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
2:  Parameter update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
4:  Direction update: According to (45), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$; update ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
5:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
V-B Modified steepest descent algorithm
In fact, alternative unconstrained optimization methods can also be employed to solve the proposed projection-based penalty problem in (31). For comparison purposes, we also develop a Modified Steepest Descent (MSD) algorithm, adapted from the proposed MCG framework. The primary distinction between the MSD and MCG algorithms lies in the choice of the descent direction. Specifically, in the MSD method, the descent direction at each iteration is directly set as the negative gradient of the objective function
${\bm{d}}_{i}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i}),$ (53)
which can be directly obtained from MCG by setting the deflection factor to zero. We herein omit the detailed implementation steps of PP-MSD-ILS since they can be obtained from the PP-MCG-ILS by simply revising step 4 of Algorithm 1.
V-C Benchmarks
We now introduce two benchmark algorithms, derived from the PP-MCG-ILS and PP-MSD-ILS frameworks, respectively. These benchmark variants adopt fixed step sizes at each iteration instead of employing ILS. To mitigate potential instability arising from overly large or small gradient magnitudes under fixed step updates, we normalize the update directions using their $l_{2}$-norm. For the Penalty-function and Projection driven Normalized Steepest Descent (PP-NSD) algorithm, the descent direction is defined as
@@ V-C Benchmarks
where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small constant to prevent division by zero. Similarly, for the penalty function and projection driven normalized CG (PP-NCG) algorithm, the descent direction is given by,
${\bm{d}}_{i+1}$ $=-({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})/(\|({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})\|_{2}+\epsilon_{d}),$ (55)
@@
The step size selection in both benchmark algorithms is
@@
with $\upsilon_{i}^{\rm b}$ provided by (51) and $\upsilon_{0}$ being a predetermined step size. Additionally, the descent function (38) may not hold when utilizing the fixed step sizes. Therefore, it is necessary to provide a sufficiently large integer $I$ as the upper bound of the number of iterations. In the same vain, due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it. To address this, we introduce an alternative penalty stopping criterion given by $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$ where $S$ represents the window size for searching the minimum penalty value over recent iterations. The detailed implementation steps of the PP-NSD algorithm are omitted here, as it can be readily constructed by modifying step 4 of the PP-NCG algorithm described in Algorithm 2.
@@
2:  Parameter update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$ or $i+1=I$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, go to step 5; Otherwise, go to next step.
4:  Direction update: According to (45), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$; update ${\bm{d}}_{i+1}$ through (55); let $i=i+1$, and go back to step 1.
5:  Penalty update: If $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})\|_{2}+\epsilon_{d})$; let $j=j+1$ and $i=i+1$, and go to step 1.
VI Extension to pure sensing
In this section, we propose a Penalty-function based Normalized Conjugate Gradient algorithm with inexact line search (P-NCG-ILS) for the PA in pure sensing scenario. We consider the power minimization under power budget constraints and CRLB constraints, i.e.,
$\arg\min\limits_{{\bm{\rho}}}$ $\ {\bm{\rho}}^{\rm T}{\bm{1}},$ (57a)
@@ VI Extension to pure sensing
We present the penalty function as follows
$\alpha_{\rm s}({\bm{\rho}})=$ ${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
@@
with the penalty function for the total power budget
${f}_{\rm s}({\bm{\rho}})=(\max\{0,{\bm{\rho}}^{\rm T}{\bm{1}}-1\})^{q}.$ (59)
Then, the objective function can be formulated as
$L_{\rm s}({\bm{\rho}})={\bm{\rho}}^{\rm T}{\bm{1}}+\mu\alpha_{\rm s}({\bm{\rho}}).$ (60)
The complete P-NCG-ILS algorithm is summarized in Algorithm 3.
@@
1:  Step size: $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\}$; while $L_{\rm s}({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L_{\rm s}({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
2:  Parameter update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
3:  Convergence condition: If $|L_{\rm s}({\bm{\rho}}_{i+1})-L_{\rm s}({\bm{\rho}}_{i})|/|L_{\rm s}({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
4:  Direction update: According to (45), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$; update ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
5:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
VII Algorithm Analysis
The proposed algorithms (e.g., PP-MCG-ILS, PP-MSD-ILS and P-NCG-ILS) can be extended and applied to a broad class of nonlinear and non-convex optimization problems. These problems are typically characterized by an affine equality constraint, if present, satisfied through a projection operator, along with nonlinear or non-convex constraints incorporated via penalty functions. The ILS mechanism is used to enhance convergence. In the following, we analyze the convergence behavior and computational complexity of the proposed algorithms.
VII-A Convergence analysis
We now provide the convergence condition of the proposed PP-MCG-ILS and PP-MSD-ILS algorithms. Before delving to the details, we define the increment of the parameter update and the objective function respectively as
$\Delta{\bm{\rho}}_{i+1}$ $\triangleq{\bm{\rho}}_{i+1}-{\bm{\rho}}_{i}=v_{i}{\bm{d}}_{i}$ (61)

### v2

@@ IV-A2 Step size optimization
where $\nabla^{2}L({\bm{\rho}}_{i})$ is the Hessian matrix of $L({\bm{\rho}})$ at ${\bm{\rho}}_{i}$. According to (35), we have ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\leq 0$. If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there exists a non-negative step size bound $\upsilon_{b}$ for given $\epsilon_{L}\in[0,1)$ such that
$\small f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}),\quad\text{for}\quad\upsilon_{i}\leq\upsilon_{b}.$ (43)
(43) always holds when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$. In contrast, we have $\upsilon_{b}=\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}$ when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}>0$.
@@
$\small L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (44)
According to (35), the term on the right-hand side of (44) is always non-positive. Thus, (44) can enable a descent function in (34). The ILS controlling factor $\epsilon_{L}$ is used to regulate the decrement of the function $L({\bm{\rho}})$ by moving ${\bm{\rho}}_{i}$ towards the direction ${\bm{d}}_{i}$ weighted by the step size $\upsilon_{i}$. When (44) is unmet for given $\upsilon_{i}$, the step size is scaled by
$\small\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i},$ (45)
where $\sigma_{\upsilon}\in(0,1)$ denotes the scaling factor and an initial step size is denoted as $\upsilon_{i}=\upsilon_{0,i}$.
Additionally, the penalty function does not inherently ensure that the individual power budget constraints at each transmitter are respected during the optimization process. As a result, negative PA values may arise in intermediate iterations. These infeasible PA values can lead to misleading CRLB evaluations, for instance, producing negative CRLBs instead of infinite ones, thereby falsely satisfying the CRLB constraints. To avoid unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is crucial to enforce non-negative PAs at each iteration step. To this end, the step size $\upsilon_{i}$ needs to satisfy
$\small 0<\upsilon_{i}\leq\upsilon_{i}^{\rm b},$ (46)
where $\upsilon_{i}^{\rm b}$ denotes an upper bound on the step size, given by
@@
with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$. Thus, the initial step size for ILS can be selected as
$\small\upsilon_{0,i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (48)
with $\epsilon_{\upsilon}\in(0,1]$. Note that there must exist at least one component ${d}_{n,i}<0$ in each iteration $i$, unless ${\bm{d}}_{i}={\bm{0}}$, due to the constraint $\sum_{n=1}^{N}{d}_{n,i}=0$. The ILS method characterized in (44)-(48) can be seen as the extension and application of the Armijo’s Rule [41] in our work.
IV-A3 The algorithm implementation
A well-defined termination condition is essential for the proposed iterative algorithm. A natural choice, inspired by the ILS condition in (44), is to monitor the absolute change in the objective function between consecutive iterations, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$, where $\epsilon$ is a predefined threshold with small value. To avoid numerical instability arising from the magnitude of the objective function itself, we apply a normalized termination condition, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$. This condition is employed in both the MCG-based and MSD-based algorithms.
We now present the complete PP-MCG-ILS. Generally, a large penalty factor may lead to numerical issues, such as an ill-conditioned Hessian matrix, which can degrade the convergence behavior of the MCG algorithm. To mitigate this, we adopt a sequential penalty strategy that gradually increases the penalty factor across successive iteration loops. Specifically, for a given penalty factor $\mu_{k}$, the MCG algorithm is executed until its termination condition is satisfied. Then, the penalty factor is updated as $\mu_{k+1}=\varphi\mu_{k}$ where $\varphi>1$ is a scaling constant, for the next loop of penalty-based iteration. This iterative refinement continues until the convergence criterion $\mu_{k}\alpha(\bm{\rho})\leq\epsilon_{\mu}$ is satisfied, where $\epsilon_{\mu}$ is a small, predefined threshold [39]. This sequential penalty strategy ensures both numerical stability and progressive enforcement of the constraint conditions. The detailed procedure of the proposed PP-MCG-ILS algorithm is summarized in Algorithm 1.
@@ IV-A3 The algorithm implementation
1:  Step size bound: Calculate $\upsilon_{i}^{\rm b}$ according to (47);
2:  Step size search: set $\upsilon_{i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$, while $L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
3:  PA update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
4:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 7; Otherwise, go to next step.
5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
6:  Direction update: ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
7:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
IV-B Modified Steepest Descent Algorithm
In fact, alternative unconstrained optimization methods can be applied to solve the proposed projection-based penalty problem in (26). In this paper, we also develop a MSD algorithm, adapted from the proposed MCG framework. The main difference between the MSD and MCG algorithms lies in how the search direction is determined. In the MSD approach, the search direction at each iteration is set to the negative gradient of the objective function, i.e.,
$\small{\bm{d}}_{i}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i}),$ (49)
which can be obtained from MCG by setting the deflection factor to zero. We herein omit the detailed implementation steps of PP-MSD-ILS since they can be obtained from the PP-MCG-ILS in Algorithm 1 by simply removing step 5 and performing the direction update in step 6 using (49).
IV-C Benchmarking Schemes
To facilitate performance comparison, we introduce two benchmark algorithms based on the PP-MCG-ILS and PP-MSD-ILS frameworks, called penalty function and projection-driven normalized SD (PP-NSD) and penalty function and projection-driven normalized CG (PP-NCG), respectively. Here, we employ the standard SD and CG methods proposed in [42], where fixed step sizes are adopted instead of ILS. To mitigate potential instability arising from overly large or small gradient magnitudes under fixed-step updates, we normalize the update directions using their $l_{2}$ norm. Specifically, for the PP-NSD algorithm, the search direction is defined as an $\epsilon_{d}$-regularized normalized gradient, given by
@@ IV-C Benchmarking Schemes
where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small positive constant introduced to prevent division by zero. Similarly, for the PP-NCG algorithm, the search direction is given by an $\epsilon_{d}$-regularized normalized conjugate gradient, i.e.,
$\small{\bm{d}}_{i+1}$ $=-({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})/(\|({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})\|_{2}+\epsilon_{d}),$ (51)
@@
The step size in both benchmark algorithms is given by
@@
with $\upsilon_{i}^{\rm b}$ given in (47) and $\upsilon_{0}$ being a predetermined step size. The maximum number of iterations is set to $I$. Due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it. To address this, we introduce an alternative penalty stopping criterion given by $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$ where $S$ represents the window size for searching the minimum penalty value. The detailed implementation steps of the PP-NSD algorithm are omitted here, as it can be easily constructed from the PP-NCG described in Algorithm 2 by removing step 4 and performing the direction update in step 5 using (50).
@@
2:  PA update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$ or $i+1=I$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, go to step 6; Otherwise, go to next step.
4:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
5:  Direction update: ${\bm{d}}_{i+1}$ through (51); let $i=i+1$, and go back to step 1.
6:  Penalty update: If $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})\|_{2}+\epsilon_{d})$; let $j=j+1$ and $i=i+1$, and go to step 1.
V Extension to Pure Sensing
In this section, we consider a pure sensing scenario with joint location and velocity estimation. The P-NCG-ILS algorithm is proposed for PA in minimizing the total power under power budget constraints and CRLB constraints.
The constrained power minimization problem is formulated as
$\small\arg\min\limits_{{\bm{\rho}}}$ $\ {\bm{\rho}}^{\rm T}{\bm{1}},$ (53a)
@@ V Extension to Pure Sensing
Analogously to the derivations of PP-MCG-ILS, we set the penalty function for violating the inequality constraints as
$\small\alpha_{\rm s}({\bm{\rho}})=$ ${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
@@
where ${f}_{\rm s}({\bm{\rho}})$ denotes the penalty sub-function for the total power budget, i.e.,
$\small{f}_{\rm s}({\bm{\rho}})=(\max\{0,{\bm{\rho}}^{\rm T}{\bm{1}}-1\})^{q}.$ (55)
Then, the unconstrained optimization problem can be formulated as
$\small\arg\min\limits_{{\bm{\rho}}}\ L_{\rm s}({\bm{\rho}})={\bm{\rho}}^{\rm T}{\bm{1}}+\mu\alpha_{\rm s}({\bm{\rho}}).$ (56)
Similarly to the PP-MCG-ILS algorithm, we also use the CG-based method to solve (56). To avoid a potential numerical instability, we consider normalized CG search direction, given by ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$, where the deflection computation method is the same as that of the PP-MCG-ILS algorithm. Additionally, ILS is used to determine the step size at each iteration. The complete P-NCG-ILS algorithm is summarized in Algorithm 3.
@@
1:  Step size bound: Calculate $\upsilon_{i}^{\rm b}$ according to (47);
2:  Step size: $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\}$; while $L_{\rm s}({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L_{\rm s}({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
3:  PA update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
4:  Convergence condition: If $|L_{\rm s}({\bm{\rho}}_{i+1})-L_{\rm s}({\bm{\rho}}_{i})|/|L_{\rm s}({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$.
6:  Direction update: ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
7:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
VI Algorithm Analysis
The proposed algorithms (e.g., PP-MCG-ILS, PP-MSD-ILS) can be extended and applied to a broad class of nonlinear and non-convex optimization problems. These problems are typically characterized by an affine equality constraint, if present, satisfied through a projection operator, along with nonlinear or non-convex constraints incorporated via penalty functions. The ILS mechanism is used to enhance convergence. For example, the proposed P-NCG-ILS algorithm for sensing-only scenario can be seen as a simple extension of the PP-MCG-ILS algorithm. In the following, we analyze the convergence behavior and computational complexity of the proposed algorithms.
VI-A Convergence Analysis
We now provide the convergence analysis of the proposed PP-MCG-ILS and PP-MSD-ILS algorithms. Before delving to the details, we define the increment of the parameter update and the objective function respectively as
$\small\Delta{\bm{\rho}}_{i+1}$ $\triangleq{\bm{\rho}}_{i+1}-{\bm{\rho}}_{i}=v_{i}{\bm{d}}_{i},$ (57)